Find C. with work shown pleaseA report announced that the median sales price of new houses sold one year was $211,000, and the mean sales price was $271,90
Answer #1

Solution :

Given that ,

\mu\bar x = 271,900

\sigma\bar x = \sigma / \sqrt n = 100,000 / \sqrt 100 = 10000

P(\bar x < 300,000) = P((\bar x - \mu \bar x ) / \sigma \bar x < (300,000 - 271,900) / 10000)

= P(z < 2.81)

Using z table

= 0.9975